OpenStudy (anonymous):

The equation below has a solution near x = 0. By replacing the left side of the equation by its linearization, find an approximate value for the solution. Note: Your answer should be in fraction form. e^x+14x=7

OpenStudy (anonymous):

i have to say it does not have a solution that near 0 because \(e^0+14\times 0=1\) not \(7\)

OpenStudy (anonymous):

i need the approximate value

OpenStudy (anonymous):

not the solution

OpenStudy (anonymous):

you can find the equation of the line tangent to the graph of \(e^x+14x\) at \((0,1)\) but i don't see how that will give you a good approximation for the equation

OpenStudy (anonymous):

derivative of \(e^x+14x-7\) is \(e^x+14\) at \(0\) the slopes is \(14\) and the equation of the tangent line at \((0,1)\) is \(y-1=14x\)

OpenStudy (anonymous):

maybe i am missing something. are you supposed to be using newton ralphson?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i was wrong anyway, at \(0\) the slope is 15

OpenStudy (anonymous):

the local linearization formula that all. well that what the chapter is about

OpenStudy (anonymous):

maybe you are supposed to be finding the zero of the line tangent to the curve of \[y=e^x+14x-7\] that is easy enough, since \(y'=e^x+14\) and at \(0\) you get 15 for the slope. the point would be \((0,-6)\) and the equation for the tangent line would be \[y+6=15x\] or \[y=15x-6\] zero is therefore \[\frac{6}{15}\]

OpenStudy (anonymous):

a rather lousy approximation, but if that is what is needed ok.

OpenStudy (anonymous):

yea i think your right